OpenStudy (anonymous):

MEDAL Which ratio is not equivalent to 24/36 ? a 3/4 b 12/18 c 4/6 d 48/72 *L_*

OpenStudy (anonymous):

It is 3/4 because the Equivalent factions are; 2/3 = 4/6 = 6/9 = 8/12 = 10/15 = 12/18 = 14/21 = 16/24 = 18/27 = 20/30 = 22/33 = 24/36 = 26/39 = 28/42 = 30/45 = 32/48 = 34/51 = 36/54 = 38/57 = 40/60

OpenStudy (anonymous):

Yeah 3/4 . Beat me to it